OpenStudy (anonymous):

Help walk me though this problem? (x^2-2x-3)/(x+1)

OpenStudy (anonymous):

factor the numerator: \(\large x^2-2x-3=(x-3)(x+1) \) now cancel common factors from numerator and denominator.

OpenStudy (anonymous):

No, I need to divide the two. My bad!

OpenStudy (anonymous):

same thing... \(\large \frac{x^2-2x-3}{x+1}=\frac{(x+1)(x-3)}{x+1}=\frac{\cancel{(x+1)}(x-3)}{\cancel{x+1}} =x-3\)